Joe and Johan walked about 20 km yesterday surveying our sampling site and collecting albedo spectra. Tris flew the quadcopters and acquired a lot of albedo footprints, and Chris, Ewa and Tom were busy collecting complimentary microbiological samples. Tris and Tom leave camp on Tuesday, so time will be going very quickly for them now, and they will be anxious to collect as much data as they can. Meanwhile, Jim is changing to sampling air for biological materials today, and is busy as a consequence.
Johan and Tris have done a sterling job repitching the personal tents, and the rest of the team learnt a few tricks about how to do this quickly as a consequence.
Field camp food fantasies continue – fresh eggs are the current craving, and Jenine, Liane and Steffi are tasked with bringing in as many as they can on Tuesday’s field crew change.